Christopher's lunch contains 60 grams of carbohydrate, 25 grams of protein, and 15 grams of fat. What percent of kilocalories (kcal) in this meal come from fat?
Answer #1

Calorie from carbo hydrates= 60× 4= 240 Cal

Calorie from protein =25×4 = 100 Cal

Calorie from fat = 15×9 = 135Cal = 0.135 Kcal

Total Calories =240+ 100+ 135 = 475Calories =0.475 K cal

Percentage of Kcal from the fat = 135 x 100 475

   = 28.42 %